### v1

# Fundamental Examples of non-compact Reeb spaces of smooth functions

Abstract.
Reeb spaces of (continuous) real-valued functions on (nice) topological spaces are the spaces whose underlying sets consist of all connected components (contours) of their level sets and seen naturally as quotient spaces of the spaces. They are ”$1$-dimensional” spaces in various nice cases. They are graphs or graphs with ends for smooth function cases with nice singularities and behaviors. These tools or objects, have been fundamental and important in theory of Morse functions and more general smooth functions and applications to geometry, since the 20th century.
We present new systematic methods to obtain Reeb spaces homeomorphic to infinite graphs (with ends) for functions on non-compact manifolds with no boundary. This note is regarded as a note on cases previously obtained by the author from a kind of general viewpoints.
Key words and phrases:
Smooth, real analytic, or real algebraic (real polynomial) functions and maps. Reeb spaces. Graphs. Reeb graphs
2020 Mathematics Subject Classification: Primary 26E05, 54C30, 57R45, 58C05.
1. Introduction.
The Reeb space $R_{c}$ of a (continuous) real-valued function $c:X\rightarrow\mathbb{R}$ on a (nice) topological space $X$ is regarded to be the space of all connected components (contours) of all preimages (level sets) $c^{-1}(q)$. In the case of differentiable functions, we add ”regular” (resp. ”critical”) before ”contour” and ”level set” if it contains no critical point (resp. some critical points) of the function.
We introduce the notions mathematically and rigorously. We can induce the equivalence relation ${\sim}_{c}$ on $X$ by the rule: the relation $x_{1}{\sim}_{c}x_{2}$ holds if and only if $x_{1}$ and $x_{2}$ are in a same contour of a same level set $c^{-1}(y)$. The quotient map $q_{c}:X\rightarrow R_{c}:=X/{{\sim}_{c}}$ and the unique continuous function $\bar{c}:R_{c}\rightarrow\mathbb{R}$ with $c=\bar{c}\circ q_{c}$ is also defined.
These tools or objects have been well-known as fundamental and important tools or objects in theory of Morse functions and more general differentiable functions with nice singularity ([14]). $R_{c}$ is, for a function $c$ which is not so wild mathematically, has the structure of a graph. For the Morse function case and Morse-Bott function case (on compact manifolds), see [2, 11] and for more general cases, see [1, 15, 16], for example. Especially, [1, Theorem 7.5], [15, Theorem 3.1], and [16, Theorems 2.1 and 2.8 and ”2 Reeb space and its graph structure”] are important as theory on topologies and graph structures of these spaces.
Hereafter, a graph means a $1$-dimensional connected CW complex such that the closure of each $1$-cell (edge) is homeomorphic to the $1$-dimensional disk $D^{1}:=\{x\in\mathbb{R}\mid x^{2}\leq 1\}$: 0-cell is a vertex. As a wider class, a graph with ends means a $1$-dimensional connected CW complex such that the closure of each $1$-cell (edge) is homeomorphic to the $1$-dimensional disk $D^{1}:=\{x\in\mathbb{R}\mid x^{2}\leq 1\}$ or a ray $\{x\geq 0\mid x\in\mathbb{R}\}$: 0-cell is a vertex as above.
In the case of the smooth manifold $X$ with no boundary, $R_{c}$ is, in a certain nice case, a graph (with ends) whose vertex set consists of all critical contours of $c$. This is the so-called Reeb graph of $c$ and in most studies including presented studies above, this definition is adopted. We can see the Reeb graph as a digraph (with ends) by giving orientation on each edge by the following rule.
  • If an edge $e$ departs from a vertex $v_{e,1}$, then it is an oriented edge departing from $v_{e,1}$.
  • If an edge $e$ enters a vertex $v_{e,2}$, then it is an oriented edge entering $v_{e,2}$.
Problem 1.
Can we construct a (nice) smooth function whose Reeb space has the structure of the Reeb graph (with ends) isomorphic to a given graph (resp. with ends)?
Hereafter, let ${\mathbb{R}}^{k}$ denote the $k$-dimensional Euclidean (real affine) space with $\mathbb{R}:={\mathbb{R}}^{1}$. Let ${\pi}_{k_{1},k_{2}}:{\mathbb{R}}^{k_{1}}\rightarrow{\mathbb{R}}^{k_{2}}$ denote the canonical projection mapping $x=(x_{1},x_{2})\in{\mathbb{R}}^{k_{1}-k_{2}}\times{\mathbb{R}}^{k_{2}}$ to $x_{1}$ with $k_{1}>k_{2}\geq 1$
This is a kind of fundamental, natural and incredibly new problems. Sharko has launched in this 2006 ([17]). This is a study on smooth functions on closed surfaces whose critical points $p$ are of the form $c(z)={\Re}z^{l}+t_{p}$ or critical points of Morse functions and whose Reeb graphs are given graphs of a certain class. Here $\mathbb{C}$ is the space of all complex numbers, $l>1$ is an integer, $t_{p}\in\mathbb{R}$, and ${\Re}s$ denotes the real part of $s\in\mathbb{C}$. This is extended to arbitrary finite graphs in [12] and functions whose critical points may not be isolated are studied. In [13], a certain class of finite graphs is important and Morse functions on closed manifolds with prescribed Reeb graphs of the class and regular contours which are spheres are reconstructed. The author has contributed to this, in, [3, 5], for example. There the author respects not only graphs, but also shapes of level sets. In addition, in [4], the author considers smooth functions on non-compact manifolds and so-called non-proper functions, explicitly, first. As another important pioneering study by the author, [6] is on real algebraic construction with prescribed Reeb graphs. The author is contributing to real algebraic construction, mainly, recently, where we ignore related knowledge, and omit precise exposition on other related papers and preprints, except [7, 8]. Later, as a kind of our main ingredients, related arguments appear in our main ingredients and we discuss the arguments in self-contained ways.
Recently the author has started to consider infinite graphs (with ends) as Reeb spaces and obtained several explicit cases. They are also constructed as, fully or densely, real algebraic or real analytic and based on the paper [5] and the preprints [7, 8]. Related preprints are [9, 10]. These preprints are important in our present paper, where we ignore related non-trivial knowledge and discuss related arguments in self-contained ways. The present paper is also a kind of our new remark on examples of [9, 10], especially [9, Theorem 1]. There real algebraic or real analytic functions or maps are presented and an answer to Problem 1 are presented as the restrictions of the canonical projections ${\pi}_{k_{1},k_{2}}$ to the zero sets of real analytic or algebraic maps on the real affine spaces. We respect these examples and present an answer from a kind of essentially generalized viewpoints.
In the next section, we present our main result and prove the new theorem (Theorem 3). This is for a generalization of the case presented in [9, Theorem 2] and in Example 1, this is explained again.
2. Our main result.
We introduce or review several notions additionally. Some already appear in the last section, where it is no problem. They are fundamental and well-known.
For a map $c:X\rightarrow Y$ and a subset $Z\subset Y$, let $c{\mid}_{Z}$ denote the restriction of $c$ to $Z$.
A proper map $c:X\rightarrow Y$ between topological spaces $X$ and $Y$ is a map whose preimage $c^{-1}(K)$ is compact for any each subset $K\subset X$. A non-proper map is one which is not proper.
For a differentiable map between differentiable manifolds, a point in the manifold of the domain is a singular point of it where the rank of the differential of the map is smaller than both the dimension of the manifold of the domain and that of the manifold of the target. The singular set of the map is defined as the set of all singular points of the map. We use ”critical” in the case where the map is the real-valued map.
Theorem 1 is essentially from [5].
Theorem 1.
Let $m\geq 2$ be an integer. Let $c_{i}:\mathbb{R}\rightarrow\mathbb{R}$ ($i=1,2$) be smooth functions such that $c_{1}(x)<c_{2}(x)$ for any $x\in\mathbb{R}$. Then consider a region $D_{c_{1},c_{2}}:=\{(x_{1},x_{2})\mid c_{1}(x_{2})<x_{1}<c_{2}(x_{2})\}$ and its closure $\overline{D_{c_{1},c_{2}}}$ in ${\mathbb{R}}^{2}$. Then we have an $m$-dimensional smooth submanifold $X_{m,c_{1},c_{2}}:=\{(x_{1},x_{2},{(y_{j})}_{j=1}^{m-1})\in\overline{D_{c_{1},c_{2}}}\times{\mathbb{R}}^{m-1}\subset{\mathbb{R}}^{m+1}\mid(x_{1}-c_{1}(x_{2}))(c_{2}(x_{2})-x_{1})-{\Sigma}_{j=1}^{m-1}{y_{j}}^{2}=0\}$ with no boundary and the restriction ${\pi}_{m+1,2}{\mid}_{X_{m,c_{1},c_{2}}}$ is a surjection onto $\overline{D_{c_{1},c_{2}}}$ whose singular set is $\{(x_{1},x_{2},{(y_{j})}_{j=1}^{m-1})\mid(x_{1},x_{2})\in\overline{D_{c_{1},c_{2}}}-D_{c_{1},c_{2}}\}=\{(x_{1},x_{2},{(0)}_{j=1}^{m-1})\mid(x_{1},x_{2})\in\overline{D_{c_{1},c_{2}}}-D_{c_{1},c_{2}}\}$.
Proof.
This is a main ingredient of [5] and we prove this in a self-contained way. It is a main ingredient to show that by implicit function theorem, this is the zero set of the smooth function and is a smooth manifold with no boundary. This is the zero set of the smooth function $(x_{1}-c_{1}(x_{2}))(c_{2}(x_{2})-x_{1})-{\Sigma}_{j=1}^{m-1}{y_{j}}^{2}$ by the definition.
In the case $(x_{1},x_{2})\in D_{c_{1},c_{2}}$ in the zero set, the value of the partial derivative of the function $(x_{1}-c_{1}(x_{2}))(c_{2}(x_{2})-x_{1})-{\Sigma}_{j=1}^{m-1}{y_{j}}^{2}$ by some $y_{j}$ is not 0.
In the case $(x_{1},x_{2})\in\overline{D_{c_{1},c_{2}}}-D_{c_{1},c_{2}}$ in the zero set, the value of the partial derivative of the function $(x_{1}-c_{1}(x_{2}))(c_{2}(x_{2})-x_{1})-{\Sigma}_{j=1}^{m-1}{y_{j}}^{2}$ by $x_{1}$ is $c_{1}(x_{2})-x_{1}$ in the case $c_{2}(x_{2})-x_{1}=0$ and $c_{2}(x_{2})-x_{1}$ in the case $x_{1}-c_{1}(x_{2})=0$. The values are both non-zero.
This means that the ranks of the differential of the function are always $1$ on the zero set. The zero set is an $m$-dimensional smooth submaifold with no boundary, We can prove the remaining easily by the situation. ∎
Theorem 2.
In Theorem 1, the following hold. We abuse the notation.
  1. (1)
    A point of $X_{m,c_{1},c_{2}}$ is a critical point of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ if and only if it is a singular point of the surjection ${\pi}_{m+1,2}{\mid}_{X_{m,c_{1},c_{2}}}$ onto $\overline{{D}_{m,c_{1},c_{2}}}$ and of the form $(c_{i}(s),s,{(0)}_{j=1}^{m-1})$, where $s$ is a critical point of the function $c_{i}$.
  2. (2)
    The function $c_{i}:\mathbb{R}\rightarrow\mathbb{R}$ ($i=1,2$) is a smooth function whose critical set is a discrete set and at least one function of $c_{1}$ and $c_{2}$ has at least one critical point. Suppose also the following.
    1. (a)
      $t\in\mathbb{R}$ goes to -$\infty$, then the values $c_{1}(t)$ and $c_{2}(t)$ go to $b_{1}\in\mathbb{R}$.
    2. (b)
      $t\in\mathbb{R}$ goes to $\infty$, then the values $c_{1}(t)$ and $c_{2}(t)$ go to $b_{2}\in\mathbb{R}$.
    3. (c)
      For $i=1,2$, each connected component $K_{c_{1},c_{2},b_{i},j}$ of the set ${{\pi}_{2,1}}^{-1}(b_{i})\bigcap\overline{D_{c_{1},c_{2}}}$ has a small compact and connected neighborhood $N_{c_{1},c_{2},b_{i},j}$ consisting of contours of the the restriction ${\pi}_{2,1}{\mid}_{\overline{D_{c_{1},c_{2}}}}$ and the complementary set of $K_{c_{1},c_{2},b_{i},j}$ in the connected component $N_{c_{1},c_{2},b_{i},j}$ consists of finitely many connected sets. In addition, these neighborhoods $N_{c_{1},c_{2},b,j}$ can be chosen to be mutually disjoint.
    Then the Reeb space of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ is regarded to be the Reeb graph.
Proof.
The statement (1) holds from our construction.
We prove the statement (2). First, the critical sets of the functions $c_{1}$ and $c_{2}$ are discrete by the assumption. From the conditions (2a, 2b) and the structures of the maps and the manifolds, the restriction of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ to the preimage of $\mathbb{R}-\{b_{1},b_{2}\}$ is proper. The critical set of the restriction of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ to the preimage of each closed subset in $\mathbb{R}-\{b_{1},b_{2}\}$ is finite, and by virtue of [15, 16], its Reeb space is regarded to be homeomorphic to a $1$-dimensional manifold or a graph which is not homeomorphic to a manifold. In addition, if at least one critical point of the function exists, then the Reeb space of the restriction of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ to the preimage of $\mathbb{R}-\{b_{1},b_{2}\}$ is regarded to be the Reeb graph. For [15, 16], related to this argument, mainly, consult [15, Theorem 3.1] and [16, Theorems 2.1 and 2.8 and ”2 Reeb space and its graph structure”]. Related to this theory, see also [1]. We consider the condition (2c). According to this, around each point in ${\bar{{\pi}_{2,1}{\mid}_{\overline{D_{c_{1},c_{2}}}}}}^{-1}(\{b_{1},b_{2}\})$, the Reeb space is homeomorphic to a $1$-dimensional manifold or a finite graph which is not homeomorphic to a manifold and regarded to be the Reeb graph (if there exists a critical point of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ there), locally. This also comes from the general theory above.
At least one of the functions $c_{1}$ and $c_{2}$ has at least one critical point, by the assumption. From this together with the arguments above, the Reeb space of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ is regarded to be the Reeb graph and a graph with ends.
This completes the proof of (2).
The following class of smooth real-valued functions on $\mathbb{R}$ is a generalized class of the so-called unimodal real-valued functions.
Definition 1.
A smooth function $c:\mathbb{R}\rightarrow\mathbb{R}$ is almost-unimodal if the following hold.
  • $c$ has the maximum $c_{\rm M}$.
  • There exists an interval $\{t\mid a_{0}\leq t\leq b_{0}\}$ and $c$ has the maximum $c_{\rm M}$ there. Furthermore, there exist a sequence $\{a_{j}\}_{j=1}^{\infty}$ such that $a_{j}<a_{j+1}$ for each $j\geq 0$ and that as $j$ goes to $\infty$, $a_{j}$ goes to $\infty$ and another sequence $\{b_{j}\}_{j=1}^{\infty}$ such that $b_{j+1}<b_{j}$ for each $j\geq 0$ and that as $j$ goes to $\infty$, $b_{j}$ goes to -$\infty$ enjoying the following.
    • Let $c_{{\rm M},a,j}$ be the maximum of $c$ on the interval $\{t\mid a_{j}\leq t\leq a_{j+1}\}$. Then $\{c_{{\rm M},a,j}\}_{j=1}^{\infty}$ is a sequence such that $c_{{\rm M},a,j+1}\leq c_{{\rm M},a,j}$ for each $j\geq 0$.
    • Let $c_{{\rm M},b,j}$ be the maximum of $c$ on the interval $\{t\mid b_{j+1}\leq t\leq b_{j}\}$. Then $\{c_{{\rm M},b,j}\}_{j=1}^{\infty}$ is a sequence such that $c_{{\rm M},b,j+1}\leq c_{{\rm M},b,j}$ for each $j\geq 0$.
Theorem 3 is our main result.
Theorem 3.
Let $c_{1}$ and $c_{2}$ be functions of Theorem 2 (2). We abuse the notation. Suppose also the following.
  1. (1)
    The critical set of $c_{1}$ is unbounded or the function $c_{1}$ is almost unimodal.
  2. (2)
    The critical set of $c_{2}$ is unbounded or the function $c_{2}$ is almost unimodal.
Then, the Reeb space of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ is regarded to be the Reeb graph and isomorphic to a finite or infinite graph.
Proof.
We choose a point $p$ of $X_{m,c_{1},c_{2}}$ mapped to a point $p_{e}$ in some edge $e$ of the Reeb graph by the quotient map and a point $p_{D}:={\pi}_{2,1}(p)\in\overline{{D}_{m,c_{1},c_{2}}}$. Let $p_{D}:=(p_{D,1},p_{D,2})$. We consider the point $p_{D,c_{2}}:=(c_{2}(p_{D,2}),p_{D,2})$. The set $\{(t,p_{D,2})\mid p_{D,1}\leq t\leq c_{2}(p_{D,2})\}$ is in $\overline{{D}_{m,c_{1},c_{2}}}$. Due to the assumption (2), we can consider some path from $p_{D,c_{2}}$ to another point $p_{s,c_{2}}$ in $\overline{{D}_{m,c_{1},c_{2}}}-{D}_{m,c_{1},c_{2}}$ and have the path departing from $p_{D}$ obtained by joining the straight segment $\{(t,p_{D,2})\mid p_{D,1}\leq t\leq c_{2}(p_{D,2})\}$ and the path in $\overline{{D}_{m,c_{1},c_{2}}}-{D}_{m,c_{1},c_{2}}$ such that the restriction of ${\pi}_{2,1}$ there is injective and that the point $(p_{s,c_{2}},{(0)}_{j=1}^{m-1})\in{{\mathbb{R}}}^{2}\times{\mathbb{R}}^{m-1}={\mathbb{R}}^{m+1}$ is a critical point of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$. Related to the assumption (1), we can argue similarly. We can see that the closure of $e$ in the Reeb graph is homeomorphic to $D^{1}$. More precisely, from the argument above, we naturally have a union of finitely many closures of edges of the Reeb graph with the union containing $p_{e}$. ∎
We explain some related examples.
Example 1.
  1. (1)
    The case $c_{1}(x):=\frac{1}{2(x^{2}+1)}{\sin}^{2}x$ and $c_{2}(x):=\frac{1}{(x^{2}+1)}$ is an explicit case of Theorem 3 with the critical set of $c_{1}$ being unbounded, $c_{2}$ being almost unimodal, and $(b_{1},b_{2})=(0,0)$. This is also [9, Theorem 1].
  2. (2)
    The case $c_{1}(x):=-\frac{1}{2(x^{2}+1)}{\sin}^{2}x$ and $c_{2}(x):=\frac{1}{(x^{2}+1)}$ seems to be an explicit case of Theorem 3 with the critical set of $c_{1}$ being unbounded, $c_{2}$ being almost unimodal, and $(b_{1},b_{2})=(0,0)$. However, the Reeb space is not homeomorphic to any graph with ends. In fact, the condition (2c) from Theorem 2 is dropped and the set ${{\pi}_{2,1}}^{-1}(b_{i})\bigcap\overline{D_{c_{1},c_{2}}}$ is connected and non-compact, and incident to infinitely many $1$-cells: the Reeb space has the structure of a $1$-dimensional cell complex which is not a CW complex.
3. Conflict of interest and Data availability.
Conflict of interest.
The author is a researcher at Osaka Central Advanced Mathematical Institute (OCAMI researcher). This is supported by MEXT Promotion of Distinctive Joint Research Center Program JPMXP0723833165. He thanks this, where he is not employed there.
Data availability.
No data other than the present article is generated.